OpenStudy (anonymous):

A ball is kicked off the ground and in the air with a starting velocity of 29 feet per second. Using the formula H(t) = -16t^2 + vt + s what is the maximum height the soccer ball reaches? 1.14.2 feet 2.13.5 feet 3.13.1 feet 4.13.8 feet

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

get t from 0 = v0-gt = 29 - (32 ft/s^2) t which is time when velocity =0 then use that t in the equation for H(t) given, setting s=0 as it starts from the ground.

OpenStudy (sidsiddhartha):

according to douglaswinslowcooper u'll get t=29/32 now h(t)=-16*(29/32)^2+(29/31)*(29)+0=13.1 feet